OpenStudy (anonymous):

(√32)^x = 4 How do I solve this?

OpenStudy (mertsj):

\[\sqrt{32}=(32)^{\frac{1}{2}}=((2)^5)^{\frac{1}{2}}=2^{\frac{5}{2}}\]

OpenStudy (mertsj):

\[4=2^2\]

OpenStudy (mertsj):

Put those facts together and solve the equation.

OpenStudy (anonymous):

Oh. Thank you!

OpenStudy (anonymous):

Wait I still don't understnad. This is what I did. I don't see what I'm doing wrong. 2^5^1/2 = 2^2 2^(5/2) = 2

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Where did the x go?

OpenStudy (anonymous):

IDK DX I don't know where it's supposed to be.

OpenStudy (mertsj):